OpenStudy (anonymous):

Pre-Algebra A student solved the eqation 7x - 5 - 5x = 15 and found x = 5. What might be the student's error?

OpenStudy (anonymous):

@EveTheAwsome Thank you for that but I need to figure out HOW the student got x = 5

OpenStudy (anonymous):

Step 1. 2x5=15 Step 2: Add 5 to both sides. 2x5+5=15+5 2x=20 Step 3: Divide both sides by 2. 2x/2=20/2 =10

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):

This is what the students mistake was 7x-5-5x=15 7x-5x=10 2x=10 x=5 He probably subtracted 5 from 15 instead of adding 5 .

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem